OpenStudy (anonymous):

Simplify. (3x-1)^2 What's the answer and how do you find it?

OpenStudy (turingtest):

\[a^2=a\cdot a\]so\[(a+b)^2=(a+b)(a+b)\]to multiply two binomials like this you use what is called the FOIL method....

OpenStudy (turingtest):

First Outer Inner Last